OpenStudy (anonymous):

How would i find the csc(2pi/3) in fraction form? i know that csc(2pi/3)=1/(sqrt3/2) but how would i simplify this?

OpenStudy (anonymous):

\[\csc \frac{ 2\pi }{ 3 }=\csc \left( \pi-\frac{ \pi }{3 } \right)=\csc \frac{ \pi }{ 3 }=\frac{ 1 }{\sin \frac{ \pi }{3 } }=\frac{ 1 }{ \frac{ \sqrt{3} }{ 2 } }\]

OpenStudy (anonymous):

so would csc just equal the reciprocal sqrt3/2?